# Investigation of magnetic ordering with spin reorientation transition and optical properties in DyCo0.5Cr0.5O3 nanomaterials

### v1

@@ Abstract
This study reports the synthesis and physical properties of polycrystalline DyCo0.5Cr0.5O3 (DCCO) nanoparticles. Analysis of the powder X-ray diffraction (XRD) pattern using Rietveld refinement showed that the compound crystallizes in an orthorhombic crystal structure with a space group Pbnm. The particle size of approximately 57 nm was confirmed through micrographs obtained from field emission scanning electron microscopy and transmission electron microscopy. The X-ray photoelectron spectroscopy (XPS) investigations identified a mixed-valence state of Co and Cr cations. Magnetic susceptibility data indicated Curie-Weiss behavior in the temperature range 140-340 K and the onset of antiferromagnetic interactions with a Néel temperature of 119 K. At 31 K, DCCO shows spin reorientation transition from (GxAyFz) to (FxCyGz). Furthermore, room-temperature magnetization measurements demonstrated the antiferromagnetic behavior with weak ferromagnetic interactions at low temperatures. Additionally, DCCO exhibited semiconducting behavior with a direct optical bandgap of 1.97 eV, indicating promise for visible-light-driven energy harvesting and catalytic applications.
@@
Perovskite oxides, Sol-gel method, Nanoparticles, Spin orientation, Semiconductor
DyCo0.5Cr0.5O3 nanoparticles are synthesized for the first time.
The average size of the nanoparticles is 57 nm.
The presence of mixed valence states of Co and Cr cations was observed.
Néel transition was observed at approximately 119 K.
UV-vis. and PL spectroscopy confirmed the semiconducting nature of DyCo0.5Cr0.5O3 nanoparticles.
1 Introduction
@@
Noticeably, rare-earth-based mixed perovskite oxides show complex magnetic structures such as half-metallicity, magnetic ordering, magnetoresistance, multiferroicity, magnetostrictive effect, etc., due to rich electronic structures and strong interaction between the 3d electrons along with strong spin-orbit coupling [7, 10, 12, 2, 13]. For instance, an investigation on YbFe0.5Cr0.5O3 and TmFe0.5Cr0.5O3 exhibits spin reorientation phenomenon due to transition of spin structure from (GxAyFz) to (FxCyGz) and a distinct magnetostrictive effect along with a negative thermal expansion that is ascribed to a magnetoelastic effect caused by repulsion between the magnetic moments of nearby transition metal ions [2]. Another investigation on DyFe0.5Cr0.5O3 shows ferroelectric and antiferromagnetic ordering below transition temperature TN1 = 261 K and strong magnetocaloric effect 11.3 J/kg.K at 4.5 tesla increased by magneto-electric interaction as a result of magnetic field and temperature-induced magnetic transition [7]. Recent investigation on RFe0.5Cr0.5O3 (R = Nd and Sm) demonstrates an enormous exchange bias (EB) effect caused by the ”pinning effect” of R on Fe/Cr spins as a result of the ferromagnetic (antiferromagnetic) interaction occurring between the R and Fe/Cr sublattices [13]. Moreover, mixed perovskite oxide shows captivating optoelectronic properties, including a favorable bandgap, significant absorbance within the visible range of the solar energy, and great photocatalytic efficacy that makes them highly appealing [3, 11, 4]. For instance, Ln2NiMnO6 (Ln = La, Eu, Dy, Lu) based solar cells show great photovoltaic performance due to a favorable bandgap very close to the Si (1.1 eV) [3]. Another investigation on La2CoMnO6 and La2CuMnO6 reveals great photocatalytic performance for the elimination of volatile organic compounds. For isopropyl alcohol, ethanol, toluene, and ethylene, 100% conversion can be achieved with La2CoMnO6 at 150 ∘C, 200 ∘C, 300 ∘C, and 350 ∘C temperatures, respectively [11]. Recent investigation on DyFe0.5Cr0.5O3 demonstrates bandgap narrowing due to two transition metals in B sites and the concurrent existence of oxygen deficiency [4]. Such unique magnetic and optical properties motivated us to explore similar types of perovskite oxide DyCo0.5Cr0.5O3 nanomaterials. Nevertheless, it is challenging to produce B-site ordered perovskite oxides due to order-disordered effects during the synthesis governed by the thermodynamic and kinetic of the order-disorder reaction [14, 15]. An investigation on R2NiMnO6 (R = La, Pr, Nd, Sm, Gd, Tb, Dy, Ho, and Y) demonstrates that perfect B site ordering can be achieved by optimizing the synthesis process [16]. On the other hand, the order-disorder effect with various distortions in the perovskites framework is responsible for the complex physical behavior of mixed perovskite oxides [10, 17]. For instance, Gd2FeCrO6 perovskite oxide exhibits a complex state of antiferromagnetic (AFM) and ferromagnetic (FM) domain at a lower temperature due to octahedral distortion [10]. Furthermore, perovskite-type cobaltites have drawn significant attention because of the thermal fluctuation of their spin states [18, 19]. The processes underlying the low-spin (LS, S = 0, te) to high-spin (HS, S = 2, te) transition of Co3+ ion and its explanation regarding Co3+ spin states have been debated for several decades [19]. Moreover, it is reliably confirmed that Co3+ ion is in its non-magnetic low-spin state up to room temperature and a high temperature of 800 K, a simple LS-HS transition is observed [18, 19]. Therefore, in the present investigation, we have prepared DyCo0.5Cr0.5O3 for the first time using a cost-effective and straightforward sol-gel method to comprehensively explore their structural, morphological, magnetic, and optical characteristics extensively, along with potential applicability as a photocatalyst.
@@ 3.2 Morphological and elemental analyses
To investigate the surface morphology of the DyCo0.5Cr0.5O3 perovskite oxide, FESEM, and TEM imaging were carried out of the sample. Figure 2(a) shows the FESEM image of the DyCo0.5Cr0.5O3 nanoparticles, and 2(b) represents the distribution of the particles in terms of their corresponding size. From the FESEM imaging, the average particle size of DCCO perovskite oxide is approximately 57 nm. Furthermore, we have determined the average crystalline size using the Scherrer formula [10] of these synthesized nanoparticles, and it is found to be around 30 nm. This size is considerably smaller than the particle size observed in the FESEM micrographs, indicating the possible occurrence of particle agglomeration. A bright field TEM image of the DCCO perovskite oxide is illustrated in figure 2(c). The range of particle size 40–100 nm obtained from the TEM image is quite comparable with the data from the FESEM imaging. Additionally, the crystal lattice planes of DCCO nanoparticles with an interplanar separation of 0.46 nm, related to its (1 0 1) plane, are obtained from the high-resolution TEM (HRTEM) image displayed in figure 2(d). In the inset of figure 2(d), the ring-like pattern of selected-area-electron-diffraction (SAED) illustrates the polycrystallinity of DCCO perovskite oxide [37].
@@ 3.3 XPS analysis
X-ray photoelectron spectroscope (XPS) was used to examine the chemical binding energies and oxidation states of each atom of the as-prepared DCCO perovskite oxide. Figures 3(a-d) represent the core level XPS spectra of the Dy-4d, Co-2p, Cr-2p, and O-1s of the synthesized sample. The XPS spectrum of Dy 4d, as shown in figure 3(a), exhibits two doublet peaks at 152.9 eV and 156.25 eV due to spin-orbit splitting, related to 4d5/2 and 4d3/2, respectively, confirming the existence of Dy3+ state in the DCCO perovskite oxide [38]. The Co 2p orbital’s spin-orbital splitting can be seen in the XPS spectrum of the Co content (Figure 3(b)) as two primary peaks near 780.1 eV and 795.85 eV, which correspond to the Co-2p3/2 and Co-2p1/2 states, respectively, and a satellite-like peak with the binding energy of 803.61 eV. Co-2p1/2’s deconvolution peak at 795.85 eV represents the Co3+ state, while Co-2p3/2’s deconvolution peak at 780.1 eV and 784.5 eV represents the Co3+ and Co2+ states, respectively [39, 40]. In DyCo0.5Cr0.5O3, the concentration of Co3+ is 67%, while the concentration of Co2+ is 33%. Figure 3(c) shows the deconvolution of Cr 2p3/2 at 575.5 eV and 577 eV, corresponding to Cr2+ and Cr3+ states and Cr 2p1/2 at 585.58 eV and 586.5 eV, corresponding to Cr2+ and Cr3+ states, respectively [41, 42]. By comparing the areas of the XPS spectra for the two valence states, [Cr2+]/[Cr3+] has been determined to be 0.76. Finally, figure 3(d) demonstrates the XPS spectrum of the O-1s1/2 orbital, the deconvolution peak near 529.02 eV is the characteristic peak of O2- ions, and the deconvolution peak near 530.58 eV describes the ionization of oxygen ions or O- species [6, 37].

### v3

# Investigation of magnetic ordering with spin reorientation transition and optical properties in Dy2CoCrO6 nanomaterials

@@ Abstract
This study reports the synthesis and physical properties of polycrystalline Dy2CoCrO6 (DCCO) nanoparticles. Analysis of the powder X-ray diffraction (XRD) pattern using Rietveld refinement showed that the compound crystallizes in an orthorhombic crystal structure with a space group Pbnm. The particle size of approximately 57 nm was confirmed through micrographs obtained from field emission scanning electron microscopy and transmission electron microscopy. The X-ray photoelectron spectroscopy (XPS) investigations identified a mixed-valence state of Co and Cr cations. Magnetic susceptibility data indicated Curie-Weiss behavior in the temperature range 140-340 K and the onset of antiferromagnetic interactions with a Néel temperature of 119 K. At 31 K, DCCO shows spin reorientation transition from (GxAyFz) to (FxCyGz). Furthermore, room-temperature magnetization measurements demonstrated the antiferromagnetic behavior with weak ferromagnetic interactions at low temperatures. Additionally, DCCO exhibited semiconducting behavior with a direct optical bandgap of 1.97 eV, indicating promise for visible-light-driven energy harvesting and catalytic applications.
@@
Perovskite oxides, Sol-gel method, Nanoparticles, Spin orientation, Semiconductor1 Introduction
@@
Noticeably, rare-earth-based mixed perovskite oxides show complex magnetic structures such as half-metallicity, magnetic ordering, magnetoresistance, multiferroicity, magnetostrictive effect, etc., due to rich electronic structures and strong interaction between the 3d electrons along with strong spin-orbit coupling [7, 10, 12, 2, 13]. For instance, an investigation on YbFe0.5Cr0.5O3 and TmFe0.5Cr0.5O3 exhibits spin reorientation phenomenon due to transition of spin structure from (GxAyFz) to (FxCyGz) and a distinct magnetostrictive effect along with a negative thermal expansion that is ascribed to a magnetoelastic effect caused by repulsion between the magnetic moments of nearby transition metal ions [2]. Another investigation on DyFe0.5Cr0.5O3 shows ferroelectric and antiferromagnetic ordering below transition temperature TN1 = 261 K and strong magnetocaloric effect 11.3 J/kg.K at 4.5 tesla increased by magneto-electric interaction as a result of magnetic field and temperature-induced magnetic transition [7]. Recent investigation on RFe0.5Cr0.5O3 (R = Nd and Sm) demonstrates an enormous exchange bias (EB) effect caused by the ”pinning effect” of R on Fe/Cr spins as a result of the ferromagnetic (antiferromagnetic) interaction occurring between the R and Fe/Cr sublattices [13]. Moreover, mixed perovskite oxide shows captivating optoelectronic properties, including a favorable bandgap, significant absorbance within the visible range of the solar energy, and great photocatalytic efficacy that makes them highly appealing [3, 11, 4]. For instance, Ln2NiMnO6 (Ln = La, Eu, Dy, Lu) based solar cells show great photovoltaic performance due to a favorable bandgap very close to the Si (1.1 eV) [3]. Another investigation on La2CoMnO6 and La2CuMnO6 reveals great photocatalytic performance for the elimination of volatile organic compounds. For isopropyl alcohol, ethanol, toluene, and ethylene, 100% conversion can be achieved with La2CoMnO6 at 150 ∘C, 200 ∘C, 300 ∘C, and 350 ∘C temperatures, respectively [11]. Recent investigation on DyFe0.5Cr0.5O3 demonstrates bandgap narrowing due to two transition metals in B sites and the concurrent existence of oxygen deficiency [4]. Such unique magnetic and optical properties motivated us to explore similar types of perovskite oxide Dy2CoCrO6 nanomaterials. Nevertheless, it is challenging to produce B-site ordered perovskite oxides due to order-disordered effects during the synthesis governed by the thermodynamic and kinetic of the order-disorder reaction [14, 15]. An investigation on R2NiMnO6 (R = La, Pr, Nd, Sm, Gd, Tb, Dy, Ho, and Y) demonstrates that perfect B site ordering can be achieved by optimizing the synthesis process [16]. On the other hand, the order-disorder effect with various distortions in the perovskites framework is responsible for the complex physical behavior of mixed perovskite oxides [10, 17]. For instance, Gd2FeCrO6 perovskite oxide exhibits a complex state of antiferromagnetic (AFM) and ferromagnetic (FM) domain at a lower temperature due to octahedral distortion [10]. Furthermore, perovskite-type cobaltites have drawn significant attention because of the thermal fluctuation of their spin states [18, 19]. The processes underlying the low-spin (LS, S = 0, te) to high-spin (HS, S = 2, te) transition of Co3+ ion and its explanation regarding Co3+ spin states have been debated for several decades [19]. Moreover, it is reliably confirmed that Co3+ ion is in its non-magnetic low-spin state up to room temperature and a high temperature of 800 K, a simple LS-HS transition is observed [18, 19]. Therefore, in the present investigation, we have prepared Dy2CoCrO6 for the first time using a cost-effective and straightforward sol-gel method to comprehensively explore their structural, morphological, magnetic, and optical characteristics extensively, along with potential applicability as a photocatalyst.
@@ 3.2 Morphological and elemental analyses
To investigate the surface morphology of the Dy2CoCrO6 perovskite oxide, FESEM, and TEM imaging were carried out of the sample. Figure 2(a) shows the FESEM image of the Dy2CoCrO6 nanoparticles, and 2(b) represents the distribution of the particles in terms of their corresponding size. From the FESEM imaging, the average particle size of DCCO perovskite oxide is approximately 57 nm. Furthermore, we have determined the average crystalline size using the Scherrer formula [10] of these synthesized nanoparticles, and it is found to be around 30 nm. This size is considerably smaller than the particle size observed in the FESEM micrographs, indicating the possible occurrence of particle agglomeration. A bright field TEM image of the DCCO perovskite oxide is illustrated in figure 2(c). The range of particle size 40–100 nm obtained from the TEM image is quite comparable with the data from the FESEM imaging. Additionally, the crystal lattice planes of DCCO nanoparticles with an interplanar separation of 0.46 nm, related to its (1 0 1) plane, are obtained from the high-resolution TEM (HRTEM) image displayed in figure 2(d). In the inset of figure 2(d), the ring-like pattern of selected-area-electron-diffraction (SAED) illustrates the polycrystallinity of DCCO perovskite oxide [37].
@@ 3.3 XPS analysis
X-ray photoelectron spectroscope (XPS) was used to examine the chemical binding energies and oxidation states of each atom of the as-prepared DCCO perovskite oxide. Figures 3(a-d) represent the core level XPS spectra of the Dy-4d, Co-2p, Cr-2p, and O-1s of the synthesized sample. The XPS spectrum of Dy 4d, as shown in figure 3(a), exhibits two doublet peaks at 152.9 eV and 156.25 eV due to spin-orbit splitting, related to 4d5/2 and 4d3/2, respectively, confirming the existence of Dy3+ state in the DCCO perovskite oxide [38]. The Co 2p orbital’s spin-orbital splitting can be seen in the XPS spectrum of the Co content (Figure 3(b)) as two primary peaks near 780.1 eV and 795.85 eV, which correspond to the Co-2p3/2 and Co-2p1/2 states, respectively, and a satellite-like peak with the binding energy of 803.61 eV. Co-2p1/2’s deconvolution peak at 795.85 eV represents the Co3+ state, while Co-2p3/2’s deconvolution peak at 780.1 eV and 784.5 eV represents the Co3+ and Co2+ states, respectively [39, 40]. In Dy2CoCrO6, the concentration of Co3+ is 67%, while the concentration of Co2+ is 33%. Figure 3(c) shows the deconvolution of Cr 2p3/2 at 575.5 eV and 577 eV, corresponding to Cr2+ and Cr3+ states and Cr 2p1/2 at 585.58 eV and 586.5 eV, corresponding to Cr2+ and Cr3+ states, respectively [41, 42]. By comparing the areas of the XPS spectra for the two valence states, [Cr2+]/[Cr3+] has been determined to be 0.76. Finally, figure 3(d) demonstrates the XPS spectrum of the O-1s1/2 orbital, the deconvolution peak near 529.02 eV is the characteristic peak of O2- ions, and the deconvolution peak near 530.58 eV describes the ionization of oxygen ions or O- species [6, 37].